# Hilbert’s Theorem 90, periodicity, and roots of Artin-Schreier polynomials

### v2

@@ 3 Periodicity and Hilbert’s additive Theorem 90
(1) $\sigma x_{i-1}-x_{i}=-z\quad\textup{for $1\leqslant i<n$, and}\quad\sigma x_{n-1}-x_{0}=1-z.$
@@
To see that these necessary equations are sufficient, choose $z\in E$ so that $\textup{Tr}_{F}^{E}(z)=1$ which is possible as the trace map $\textup{Tr}_{F}^{E}\colon E\to F$ is surjective. Next, define $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where the empty sum $x_{0}=\sum_{j=1}^{0}\sigma^{j}z$ is 00 by convention. Then the $n$ equations in (1) hold, and $\sigma x-x$ equals $y$, as claimed. ∎
@@
Swapping the order of summation, swapping $i\leftrightarrow j$, and noting $\textup{Tr}^{E}_{F}(zy)=\sum_{i=0}^{n-1}\sigma^{i}(zy)$ gives
@@
(a) First, we show that the fixed subfield $K$ of $\sigma^{n_{p}}$ contains an element $z$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Since $n_{p^{\prime}}=n/n_{p}\neq 0$ in $F$ and the map $\textup{Tr}^{K}_{F}\colon K\to F$ is surjective, there exists an element $z\in K$ such that $\textup{Tr}^{K}_{F}(z)=(n_{p^{\prime}})^{-1}$. Therefore $\textup{Tr}^{E}_{K}(z)=|E:K|z=n_{p^{\prime}}z$ and hence $\textup{Tr}^{E}_{F}(z)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(z))=n_{p^{\prime}}\textup{Tr}^{K}_{F}(z)=1$. Next, suppose that $L=F(z)$ and $\textup{Tr}^{E}_{F}(z)=1$. Then $1=\textup{Tr}^{L}_{F}(\textup{Tr}^{E}_{L}(z))=|E:L|\textup{Tr}^{L}_{F}(z)$. Therefore $p\nmid|E:L|$, so $n_{p}=|K:F|$ divides $e=|L:F|$ as claimed. In summary, the minimal polynomial over $F$ of an element $z$ with $\textup{Tr}^{E}_{F}(z)=1$ has degree a multiple of $n_{p}$, and degree precisely $n_{p}$ if $z$ lies in $K\setminus K_{0}$ where $K_{0}$ is the unique (maximal) subfield of $K$ with $F\leqslant K_{0}<K$ and $|K:K_{0}|=p$.
(b) Let $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where $\textup{Tr}^{E}_{F}(z)=1$. Let $L=F(z)$ and $e=|L:F|$. Hence $x_{e}=\textup{Tr}^{L}_{F}(z)\neq 0$ by part (b) and $\sigma^{j}z=z$ implies $e\mid j$. Suppose that $x_{k}=0$ where $0\leqslant k<pe$ and write $k=ie+j$ where $0\leqslant i<p$ and $0\leqslant j<e$. The equation $x_{ie+j}=ix_{e}+x_{j}$ shows that $x_{0}=x_{pe}=0$. Thus $x_{k}=ix_{e}+x_{j}=0$ and $x_{e}\in F$ gives
@@
Whence $\sigma^{j}z=x_{j+1}-x_{j}=z$. This implies that $e\mid j$ and hence $j=0$. However, $0=x_{k}=x_{ie}=ix_{e}$ shows that $0=i=j=k$. Therefore $x_{\ell}\neq 0$ holds for $0<\ell<pe$, and the sequence $x_{0},x_{1},\dots$ is periodic, with period precisely $pe$. ∎
@@ 4 Formulas for the roots of Artin-Schreier polynomials
In this section we assume that $F$ has prime characteristic $p$. Set $q=p^{f}$. Two examples with $\textup{Gal}(E/F)=\langle\sigma\rangle$ of order $n$ and $\sigma(x)=x^{q}$ for $x\in E$ are the finite fields $E={\mathbb{F}}_{q^{n}}$, $F={\mathbb{F}}_{q}$, and the function fields $E={\mathbb{F}}_{q^{n}}(u)$, $F={\mathbb{F}}_{q}(u)$. We shall give a number of different applications of Theorem 1.2 depending on the value of $n_{p}$. Observe first that Theorem 1.1 gives a polynomial formula for $x$ when $\sigma(x)=x^{q}$ as noted below.
@@
Suppose that $\textup{char}(F)=p$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$, then $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{q^{j}}y^{q^{i}}$ satisfies $x^{q}-x-y=0$.
@@
Suppose that ${\mathbb{F}}_{q}\leqslant F$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$. If $y\in E$, $\textup{Tr}^{E}_{F}(y)=0$ and $n_{p}=1$, then $t^{q}-t-y=\prod_{u\in{\mathbb{F}}_{q}}(t-x-u)$ where $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$.
@@
When $\textup{char}(F)=p\nmid n$, Corollary 4.2 implies that all the roots of $t^{q}-t-y$ lie in the ${\mathbb{F}}_{p}$-subspace $\langle y,y^{q},\dots,y^{q^{n-1}}\rangle$ of ${\mathbb{F}}_{q^{n}}$. Since $\sum_{i=0}^{n-1}y^{q^{i}}=0$, the roots of $t^{q}-t-y$ lie in the (potentially larger) ${\mathbb{F}}_{q}$-subspace $\{w\in E\mid\textup{Tr}^{E}_{F}(w)=0\}$ of $E$ of codimension 1.
@@
It follows from $n_{p}=p$, that $n_{p^{\prime}}=n/p$ is coprime to $p$. Let $K$ be the fixed field of $\sigma^{p}$. Then $K$ contains an element $\zeta$ satisfying $\zeta^{p}-\zeta^{p-1}+1=0$ and $\textup{Tr}^{L}_{F}(\zeta)=1$ by Lemma 2.2. Therefore $\textup{Tr}^{E}_{F}(\zeta)=n/p$ and $\textup{Tr}^{E}_{F}(z)=1$ where $z=(n/p)^{-1}\zeta$. By Theorem 1.1, $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, $\sigma(\zeta)=\zeta^{p}$ implies that $\zeta=\sigma^{p}(\zeta)=\zeta^{p^{p}}$, so that $\zeta^{p^{p}-1}=1$ as $\zeta\neq 1$. ∎
We shall now apply Corollary 4.1 in the case that $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$, so that $E/F$ is cyclic and $\textup{Gal}(E/F)=\langle\sigma\rangle$ has order $n$ where $\sigma(x)=x^{p}$ is the $p$th power map.
Let $p,r$ be distinct primes. Denote by $\textup{ord}_{r}(p)$ the smallest positive integer $e$ such that $p^{e}\equiv 1\pmod{r}$ and note that $e\mid(r-1)$. The cyclotomic polynomial $\Phi_{r}(t)=\sum_{i=0}^{r-1}t^{i}$ of degree $r-1$ factors over ${\mathbb{F}}_{p}$ as a product of $(r-1)/e$ irreducible polynomials each of degree $e=\textup{ord}_{r}(p)$ by [6, Theorem 2.47].
@@
First, $r$ divides $p^{d}-1$ if and only if $e\mid d$. As $e\mid n$, all primitive $r$th roots of unity lie in $K$, and $F\leqslant K\leqslant E$ holds. The cyclotomic polynomial $\Phi_{r}(t)$ factors over $F$ as a product of $(r-1)/e$ irreducible polynomials each of degree $e$, so the coefficient of $t^{e-1}$ must be nonzero in at least one irreducible. Hence some $r$th root of unity $\zeta\in K$ has $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. Thus $\textup{Tr}^{E}_{K}(\zeta)=\frac{n}{e}\zeta$ and $\textup{Tr}^{E}_{F}(\zeta)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(\zeta))=\frac{n}{e}\textup{Tr}^{K}_{F}(\zeta)=\frac{n\tau}{e}\neq 0$. Therefore $z=(\frac{n\tau}{e})^{-1}\zeta\in K^{\times}$ satisfies $\textup{Tr}^{E}_{F}(z)=1$. The formula for $x=R(y,z)$ follows from Theorem 1.1 and the periodicity $ep$ follows from Theorem 1.2(b). ∎
Corollary 4.6.
Let $p$ be a prime where $p\equiv 2\pmod{3}$. Let $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$ where $n$ is even and let $\omega\in E$ satisfy $\omega^{2}+\omega+1=0$. If $p\nmid(n/2)$, and $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=0$, then the sum $x=(\frac{n}{2})^{-1}\sum_{i=0}^{n-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$ where $r_{i}\in\{0,1\}$ and $i\equiv r_{i}\pmod{2}$. Finally, the sequence $x_{0},x_{1},\dots$ is periodic of period $2p$ where $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$ and $r_{i}=i\mod 2\in\{0,1\}$.
@@
Since $n$ is even and $p^{n}\equiv 1\pmod{3}$, $E$ contains a primitive cube root $\omega$ of 1. The condition $p\equiv 2\pmod{3}$ is equivalent to $\textup{ord}_{3}(p)=2$. Thus we may use Lemma 4.5 with $r=3$ and $\zeta=\omega$. However, $\omega^{p^{k}}$ is $\omega$ if $k$ is even, and $\omega^{2}$ if $k$ is odd. If $i=0,1,2,3,4,5,6$, then $s_{i}=\sum_{k=0}^{i-1}\omega^{p^{k}}$ equals 00, $\omega$, $\omega+\omega^{2}=-1$, $-1+\omega$, $-2$, $-2+\omega$, $-3$, etc. Therefore $s_{i}=-(i-r_{i})/2+r_{i}\omega$ where $r_{i}\equiv i\pmod{2}\in\{0,1\}$ and $r_{i}\in\{0,1\}$. The values of $x_{i}=(\frac{n}{2})^{-1}\left(\frac{i-r_{i}}{2}-r_{i}\omega\right)$ are periodic of period $2p$ by Theorem 1.2(b). Finally, $\frac{i-r_{i}}{2}=\lfloor\frac{i}{2}\rfloor$. ∎
The following result gives different quadratic ‘formulas’ over the finite field ${\mathbb{F}}_{2^{n}}$ that depend on the 2-part $n_{2}$ of $n$ and the choice of an element $z\in{\mathbb{F}}_{2^{n}}$ with absolute trace 1. The condition $n_{2}=2^{k}$ is equivalent to $n\equiv 2^{k}\pmod{2^{k+1}}$.
@@
Let $y\in{\mathbb{F}}_{2^{n}}$ where $\sum_{i=0}^{n-1}y^{2^{i}}=0$. Then $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$ is a root of $t^{2}+t+y$ where $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ and $z\in{\mathbb{F}}_{2^{n}}$ has absolute trace $\sum_{i=0}^{n-1}z^{2^{i}}=1$ and has degree $n_{2}$. Write $z=(n/n_{2})^{-1}\zeta$. Table 4.7 lists choices for $\zeta$ and the minimal polynomial $m_{\zeta/{\mathbb{F}}_{2}}(t)$ of $\zeta$, as a function of the value of $n_{2}$.
$n_{2}$ $1$ $2$ $4$ $8$ $16$cdots
$\zeta$ $1$ $\omega$ $\alpha$ $\beta$ $\gamma$cdots
$m_{\zeta/{\mathbb{F}}_{2}}(t)$ $t\kern-2.0pt+\kern-2.0pt1$ $t^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{4}\kern-2.0pt+\kern-2.0ptt^{3}\kern-2.0pt+\kern-2.0pt1$ $t^{8}\kern-2.0pt+\kern-2.0ptt^{7}\kern-2.0pt+\kern-2.0ptt^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{16}\kern-2.0pt+\kern-2.0ptt^{15}\kern-2.0pt+\kern-2.0ptt^{8}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$cdots

### v3

@@ 3 Periodicity and Hilbert’s additive Theorem 90
(1) $\sigma x_{i-1}-x_{i}=-z\qquad\textup{for $1\leqslant i<n$, and}\qquad\sigma x_{n-1}-x_{0}=1-z.$
@@
To see that these necessary equations are sufficient, choose $z\in E$ so that $\textup{Tr}_{F}^{E}(z)=1$ which is possible as the trace map $\textup{Tr}_{F}^{E}\colon E\to F$ is surjective. Next, define $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where the empty sum $x_{0}=\sum_{j=1}^{0}\sigma^{j}z$ is 0 by convention. Then the $n$ equations in (1) hold, and $\sigma x-x$ equals $y$, as claimed. ∎
@@
By swapping the order of summation, by swapping $i\leftrightarrow j$, and by using the formula $\textup{Tr}^{E}_{F}(zy)=\sum_{i=0}^{n-1}\sigma^{i}(zy)$, the previous display may be written
@@
(a) First, we show that the fixed subfield $K$ of $\sigma^{n_{p}}$ contains an element $z$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Let $n_{p^{\prime}}\coloneq n/n_{p}$ be the $p^{\prime}$-part of $n$. Since $n_{p^{\prime}}\neq 0$ in $F$ and the map $\textup{Tr}^{K}_{F}\colon K\to F$ is surjective, there exists an element $z\in K$ such that $\textup{Tr}^{K}_{F}(z)=(n_{p^{\prime}})^{-1}$. Therefore $\textup{Tr}^{E}_{K}(z)=|E:K|z=n_{p^{\prime}}z$ and hence $\textup{Tr}^{E}_{F}(z)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(z))=n_{p^{\prime}}\textup{Tr}^{K}_{F}(z)=1$. Next, suppose that $z\not\in K$ and $\textup{Tr}^{E}_{F}(z)=1$. Let $L=F(z)$. Then $1=\textup{Tr}^{L}_{F}(\textup{Tr}^{E}_{L}(z))=|E:L|\textup{Tr}^{L}_{F}(z)$. Therefore $p\nmid|E:L|$, so $n_{p}=|K:F|$ divides $e=|L:F|$ as claimed. In summary, the minimal polynomial over $F$ of an element $z$ with $\textup{Tr}^{E}_{F}(z)=1$ has degree a multiple of $n_{p}$, and degree precisely $n_{p}$ if $z$ lies in $K\setminus K_{0}$ where $K_{0}$ is the unique (maximal) subfield of $K$ with $F\leqslant K_{0}<K$ and $|K:K_{0}|=p$.
(b) Let $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where $\textup{Tr}^{E}_{F}(z)=1$. Let $L=F(z)$ and $e=|L:F|$. Hence $x_{e}=\textup{Tr}^{L}_{F}(z)\neq 0$ by part (b) and $\sigma^{j}z=z$ implies $e\mid j$. Suppose that $x_{k}=0$ where $0\leqslant k<pe$ and write $k=ie+j$ where $0\leqslant i<p$ and $0\leqslant j<e$. The equation $x_{\ell e+j}=\ell x_{e}+x_{j}$ shows that $x_{0}=x_{pe}=0$. Thus $x_{k}=ix_{e}+x_{j}=0$ and $x_{e}\in F$ gives
@@
Whence $\sigma^{j}z=x_{j+1}-x_{j}=z$. This implies that $e\mid j$ and hence $j=0$. However, $0=x_{k}=x_{ie}=ix_{e}$ and $x_{e}\neq 0$ shows that $0=i=j=k$. Therefore $x_{\ell}\neq 0$ holds for $0<\ell<pe$, and the sequence $x_{0},x_{1},\dots$ is periodic, with period precisely $pe$. ∎
@@ 4 Formulas for the roots of Artin-Schreier polynomials
In this section we assume that $F$ has prime characteristic $p$. Set $q=p^{f}$. Two examples with $\textup{Gal}(E/F)=\langle\sigma\rangle$ cyclic of order $n$ and $\sigma(x)=x^{q}$ for $x\in E$, are the finite fields $E={\mathbb{F}}_{q^{n}}$, $F={\mathbb{F}}_{q}$, and the function fields $E={\mathbb{F}}_{q^{n}}(u)$, $F={\mathbb{F}}_{q}(u)$. We shall give a number of different applications of Theorem 1.2 depending on the value of $n_{p}$. Observe that an immediate corollary of Theorem 1.1 is a polynomial formula for a root of $t^{q}-t-y$.
@@
Suppose that $\textup{char}(F)=p$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$ for $x\in E$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{q^{j}}y^{q^{i}}$ satisfies $x^{q}-x-y=0$.
@@
Suppose that ${\mathbb{F}}_{q}\leqslant F$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$ for $x\in E$. If $y\in E$, $\textup{Tr}^{E}_{F}(y)=0$ and $n_{p}=1$, then the factorization $t^{q}-t-y=\prod_{u\in{\mathbb{F}}_{q}}(t-x-u)$ holds where $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$.
@@
When $\textup{char}(F)=p\nmid n$, Corollary 4.2 implies that all the roots of $t^{q}-t-y$ lie in the ${\mathbb{F}}_{p}$-subspace ${\mathbb{F}}_{q}+\{\sum_{i=0}^{n-1}\lambda_{i}y^{q^{i}}\mid\lambda_{1},\dots,\lambda_{n-1}\in{\mathbb{F}}_{p}\}$ of ${\mathbb{F}}_{q^{n}}$.
@@
It follows from $n_{p}=p$, that $n/p$ is coprime to $p$. Let $K$ be the fixed field of $\sigma^{p}$. Then $|K:F|=p$ and $K=F(\zeta)$ where $\zeta$ satisfies $\zeta^{p}-\zeta^{p-1}+1=0$ and $\textup{Tr}^{K}_{F}(\zeta)=1$ by Lemma 2.2. Therefore $\textup{Tr}^{E}_{F}(\zeta)=n/p$ and $\textup{Tr}^{E}_{F}(z)=1$ where $z=(n/p)^{-1}\zeta$. By Theorem 1.1, $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, $\sigma(\zeta)=\zeta^{p}$ implies that $\zeta=\sigma^{p}(\zeta)=\zeta^{p^{p}}$, so that $\zeta^{p^{p}-1}=1$. ∎
We shall now apply Corollary 4.1 in the case that $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$. In this case $E/F$ is cyclic and $\textup{Gal}(E/F)=\langle\sigma\rangle$ has order $n$ where $\sigma(x)=x^{p}$ is the $p$th power map.
Let $p,r$ be distinct primes. Denote by $\textup{ord}_{r}(p)$ the smallest positive integer $e$ such that $p^{e}\equiv 1\pmod{r}$ and note that $e\mid(r-1)$. The cyclotomic polynomial $\Phi_{r}(t)=\sum_{i=0}^{r-1}t^{i}$ of degree r1r-1 factors over ${\mathbb{F}}_{p}$ as a product of (r1)/e(r-1)/e irreducible polynomials each of degree $e=\textup{ord}_{r}(p)$ by [7, Theorem 2.47].
@@
Definition 4.5.
Call a non-zero polynomial $f(t)=\sum_{i=0}^{d}f_{i}t^{i}$ big if $d=0$ or $f_{d}f_{d-1}\neq 0$, and small if $f_{d}\neq 0$ and $f_{d-1}=0$. As a small polynomial times a small polynomial is small, and a big polynomial times a small polynomial is big, we have the following lemma.
Lemma 4.6.
If a big polynomial factors, then at least one of its factors is big.
@@
First, $r$ divides $p^{d}-1$ if and only if $e\mid d$. As $e\mid n$, all primitive $r$th roots of unity lie in $K={\mathbb{F}}_{p^{e}}$, and $F\leqslant K\leqslant E$ holds. The cyclotomic polynomial $\Phi_{r}(t)=\sum_{i=0}^{r-1}t^{i}$ is big, and $\Phi_{r}(t)$ factors over $F$ as a product of irreducible polynomials, each of degree $e$, at least one of which is big by Lemma 4.6. Hence some $r$th root of unity $\zeta\in K$ has $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. Thus $\textup{Tr}^{E}_{K}(\zeta)=\frac{n}{e}\zeta$ and $\textup{Tr}^{E}_{F}(\zeta)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(\zeta))=\frac{n}{e}\textup{Tr}^{K}_{F}(\zeta)=\frac{n\tau}{e}\neq 0$. Therefore $z=(\frac{n\tau}{e})^{-1}\zeta\in K^{\times}$ satisfies $\textup{Tr}^{E}_{F}(z)=1$. The formula for $x=R(y,z)$ follows from Theorem 1.1 and the periodicity $ep$ and $n_{p}\mid e$ follows from Theorem 1.2(b). ∎
Lemma 4.8.
Let $p$ be a prime where $p\equiv 2\pmod{3}$. Let $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$ where $n$ is even and let $\omega\in E$ satisfy $\omega^{2}+\omega+1=0$. If $p\nmid(n/2)$, and $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=0$, then the sum $x=(\frac{n}{2})^{-1}\sum_{i=0}^{n-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$ where $r_{i}=i-2\lfloor i/2\rfloor$. Finally, the sequence $x_{0},x_{1},\dots$ is periodic of period $2p$ where $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$.
@@
Since $n$ is even and $p^{n}\equiv 1\pmod{3}$, $E$ contains a primitive cube root $\omega$ of 1. The condition $p\equiv 2\pmod{3}$ is equivalent to $e=\textup{ord}_{3}(p)=2$. Thus we may use Lemma 4.7 with $r=3$, $\zeta=\omega$ and $\tau=-1$. However, $\omega^{p^{k}}$ equals $\omega$ if $k$ is even, and $\omega^{2}$ if $k$ is odd. If $i=0,1,2,3,4,5,6$, then $s_{i}=\sum_{k=0}^{i-1}\omega^{p^{k}}$ equals 0, $\omega$, $\omega+\omega^{2}=-1$, $1+\omega$, 2-2, $2+\omega$, 3-3, etc. Therefore $s_{i}=-(i-r_{i})/2+r_{i}\omega$ where $r_{i}=i-2\lfloor i/2\rfloor\in\{0,1\}$. Consequently, the values of $x_{i}=(\frac{n}{2})^{-1}\left(\frac{i-r_{i}}{2}-r_{i}\omega\right)$ are periodic of period $2p$ by Theorem 1.2(b). ∎
The following result gives different quadratic ‘formulas’ over the finite field ${\mathbb{F}}_{2^{n}}$ that depend on $n_{2}$ and the choice of an element $z\in{\mathbb{F}}_{2^{n}}$ with absolute trace 1. The condition $n_{2}=2^{k}$ is equivalent to $n\equiv 2^{k}\pmod{2^{k+1}}$. The case $k=0$ is handled by Corollary 4.2.
@@
Let $y\in{\mathbb{F}}_{2^{n}}$ where $\sum_{i=0}^{n-1}y^{2^{i}}=0$. Then $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$ is a root of $t^{2}+t+y$ where $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ and $z\in{\mathbb{F}}_{2^{n}}$ has absolute trace $\sum_{i=0}^{n-1}z^{2^{i}}=1$ and has degree $n_{2}$ and the sequence $(x_{i})$ has period $2n_{2}$. Table 1 lists choices for $z$ and the minimal polynomial $m_{z/{\mathbb{F}}_{2}}(t)$ of $z$ over ${\mathbb{F}}_{2}$, as a function of the $2$-part $n_{2}$ of $n$.
Table 1: Choices for $z$ and its minimal polynomial $m_{z/{\mathbb{F}}_{2}}(t)$ over ${\mathbb{F}}_{2}$ of degree $n_{2}$
$n_{2}$ 22 44 88 $16$ $32$
$z$ $\omega$ $\alpha$ $\beta$ $\gamma$ $\delta$
$m_{z/{\mathbb{F}}_{2}}(t)$ $t^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{4}\kern-2.0pt+\kern-2.0ptt^{3}\kern-2.0pt+\kern-2.0pt1$ $t^{8}\kern-2.0pt+\kern-2.0ptt^{7}\kern-2.0pt+\kern-2.0ptt^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{16}\kern-2.0pt+\kern-2.0ptt^{15}\kern-2.0pt+\kern-2.0ptt^{8}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{32}\kern-2.0pt+\kern-2.0ptt^{31}\kern-2.0pt+\kern-2.0ptt^{3}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$